the cost of a telephone call is 5 pesos for the first three minutes and an additional of 2 pesos for each additional minute or fractional part thereof.
find the cost of all call lasting .5 min, 3 minutes, and 5 minutes

Answer by tommyt3rd(5050)

c(t)=5+2(t-3), t>3
a) c(0.5)=5
b) c(3)=5
c) c(5)=5+2(5-3)=5+2(2)=9